THE MILLIMETER- AND SUBMILLIMETER-WAVE SPECTRUM OF CYANOFORMAMIDE (NCCONH_{2})

\underline{M. WINNEWISSER}, R. A. H. BUTLER and F. C. DE LUCIA, Department of Physics, The Ohio State University, Columbus, OH 43210; K. V. L. N. SASTRY, Department of Physics, University of New Brunswick, Fredericton, New Brunswick E3B 5A3, Canada.

If water (H_{2}O) and cyanogen (NCCN) are made to react in the presence of a catalyst, the product is cyanoformamide (NCCONH_{2}). Using this procedure Christiansen^{a} in the GIESSEN Laboratory has synthesized this molecule and has measured for the first time the microwave spectrum between 18 and 40 GHz with a Hewlett-Packard spectrometer. Employing a more convenient synthesis for NCCONH_{2}, we were able to extend the measurements into the millimeterand submillimeter region using the fast-scan spectrometer (FASSST). The chemical synthesis, the measurements, the laboratory data and their analysis will be discussed. They should provide the information necessary to search for this species in space.

aJ. J. Christiansen, to be published